OpenStudy (anonymous):

6 is what percent of 75?

OpenStudy (anonymous):

This is like saying 75 * (x/100) = 6 Solve for x

OpenStudy (anonymous):

so we have to multiply both sides by 100

OpenStudy (anonymous):

formula is is/of = %/100 so put values and evaluate 6/75 = %/100 8

OpenStudy (anonymous):

we get 75x = 600 now divide both sides by 75 x = 600/75

OpenStudy (anonymous):

this is equal to 8%

OpenStudy (anonymous):

yeah, sorry it's 8% not 8